OpenStudy (anonymous):

If you throw a dice 6 times, what's the chance that you'd get a six on

OpenStudy (anonymous):

1/6^6

OpenStudy (amistre64):

on .... what?

OpenStudy (anonymous):

a: exactly one of the throws. b: one or more of the throws.

OpenStudy (anonymous):

1/6

OpenStudy (amistre64):

\[C(6,1)\frac{1}{6}^1\frac{5}{6}^5\]

OpenStudy (amistre64):

and the other is: \[1-C(6,0)\frac{1}{6}^0\frac{5}{6}^6\]

OpenStudy (amistre64):

C(6,1) is notation for C(n,r), or \(^nC_r\) or some other form of it

OpenStudy (anonymous):

ty

OpenStudy (amistre64):

yep